Problem 4

Consider the following economy:

Consumption ExpenditureDITU UN Marginal Propensity to Import (nx) 0.04 (a) Calculate the equilibrium level of income. (0.5 mark) (b) Calculate auton

446,832 million

Planned Investment Expenditure

346,877 million

Government Expenditure

446,832 million

Exports

402,443 million

Imports

388,374 million

Marginal Propensity to Save

0.3

Marginal Tax Rate

0.32

Autonomous Taxes

301,240 million

Marginal Propensity to Import (nx)

0.04

(a) Calculate the equilibrium level of income.  (0.5 mark)

(b) Calculate autonomous consumption.  (0.5 mark)

(c) Calculate autonomous net exports.  (0.5 mark)

(d) Calculate autonomous planned expenditures.  (0.5 mark)

(e) Calculate the marginal leakage rate.  (0.5 mark)

(f) Assume that the natural rate of output for this economy is estimated as $1,200,000 million.

(i) Is this economy facing a recessionary or inflationary gap?  (0.5 mark)

(ii) Illustrate the gap you identified in part (i) above using both the AS-AD Model and the Aggregate Expenditure Model.  (1 mark)

(iii) Calculate the output ratio for this economy (0.5 mark)

(iv) If the government wishes to move the economy to its natural rate, will it need to increase or decrease spending? Calculate by how much it will need to change its spending.  (1 mark)

(v) Consider the policy action undertaken in part (iv) above and illustrate the impact on the money market.  (1 mark)

(vi) Given the impact on the money market determined in part (v) above, explain how this could affect the exchange rate market.  (1.5 marks)

(vii) Explain the policy action the government could undertake if it decides that it wants to move the economy to its natural rate but doesn’t want the action to affect its budget position.  (2 marks)
Answer #1

Solution-

a) At the equilibrium level,

Y=C+I+G+NX, where NX= exports-imports

Y= 446832+346877+446832+402443-388374

Y= $1,254,610 million

b) C= Ca+mpc×Yd

Here Ca= Autonomous level of Consumption, Mpc is marginal propensity to consume and Yd is disposable income.

Yd=Y-Ta-tY

Yd= 1254610-301,240-0.32×1,254,610

Yd= 551,894.8

So, Ca= C-mpc×Yd

Ca= 446832-0.7×551894.8

Ca= $60505.64

c) Net exports= Autonomous net exports-mpi×Y

NXa=NX+nxY

NXa= 14069+0.04×1254610=$64253.4

d) Expenditure multiplier= 1/1-c+ct+nx(m)

1/0.3+0.7×0.32+.04=1/0.564=1.77305

So, planned expenditures= Autonomous planned expenditures+expenditure multiplier ×Y

793709=autonomous planned expenditures+ 1.77305×1254610

A.P.E= 793709-1.77305×1254610=$ -1430777

According to the Chegg guidelines we are required to solve only till d subparts, for the rest of the parts post separately.